# A Survey on Text-guided 3D Visual Grounding: Elements, Recent Advances, and Future Directions

### v1

@@ 3.1 Benchmark Datasets
Achieving optimal performance in the T-3DVG task involving 3D vision and language heavily relies on the availability of extensive datasets. However, the process of acquiring and annotating volumetric data presents formidable hurdles, resulting in smaller collections of RGB-D datasets [72, 73, 74, 75] for three-dimensional scenarios compared to their two-dimensional counterparts. Consequently, researchers often resort to employing synthetic data to supplement the dearth of authentic data from the real world [76, 77]. To tackle this challenge, Dai et al. pioneered the creation of the ScanNet dataset [32], an extensively annotated compilation of three-dimensional indoor scans derived from real-world settings, captured through a proprietary RGB-D capture mechanism.
Comprising 2.5 million perspectives derived from 1513 scans, drawn from 707 unique environments, the ScanNet dataset surpasses other well-known datasets [78, 79, 73, 80, 75] like NYUv2 [78], SUN3D [73], and SceneNN [75] in terms of scale. Nevertheless, it is essential to recognize that ScanNet primarily caters to tasks such as categorizing 3D objects, labeling semantic voxels, and retrieving 3D objects with annotations at the instance level within a 3D point cloud. Consequently, it may not be directly applicable to tasks like 3D visual grounding and 3D dense captioning. To address this constraint, novel datasets such as ScanRefer [7], Nr3D and Sr3D [8] have emerged, specifically tailored for 3D visual grounding tasks, building upon the foundation laid by ScanNet. Both ScanRefer and Nr3D/Sr3D curate subsets of scenes from ScanNet, accompanied by additional manually crafted descriptions for the objects within each scene. Moreover, they provide accessible online platforms for data visualization, rendering them invaluable assets for the development and assessment of models in the realm of 3D visual grounding tasks.
@@
Figure 9: Demonstration of the statistics of 18 general categories, consisting of 256 sub-categories, represented within a 3D indoor environment, predominantly featuring furniture items like chairs, tables, and cabinets. The terms ”Maximum,” ”Average,” and ”Minimum” signify the highest, mean, and lowest counts, respectively, for a specific type of object present in a scene. For instance, a scene may contain as many as 32 chairs.
@@
Figure 10: The taxonomy of T-3DVG methods based on different methods. The fully-supervised method and weakly-supervised method will be introduced in Sections 4.1 and 4.2. The methods of using other modality input and using large models will be described in Sections 4.3 and 4.4. In fully-supervised methods, we present the development of T-3DVG from two perspectives, including two-stage methods and one-stage methods.
3.2 Evaluation Metrics
In the realm of 3D grounding assessment, there’s a necessity to comparing predicted locations, typically depicted as bounding boxes, against ground truth locations of objects within testing samples. Acc@$K$IoU [7] stands as the predominant metric in 3D visual grounding, quantifying the proportion of positive predictions surpassing a designated intersection over union (IoU) threshold, typically denoted as $K$ and commonly set at 0.25 or 0.5. It’s notable that certain datasets present varying evaluation configurations. For instance, in datasets like Nr3D and Sr3D, predefined object proposals are provided, with the metric focusing on the precision of selecting the target bounding box from these proposals. Additionally, various datasets assess performance across diverse scenarios. ScanRefer [7] partitions datasets into unique/multiple/overall splits, while ReferIt3D [8] categorizes them into easy/hard/view-dependent/view-independent/overall divisions. Specifically, in the ScanRefer dataset, the “unique” split includes instances where only a single distinct object from a specific category fits the description, while the “multiple” split involves cases with several objects from the same category, creating ambiguity. For example, a scene with only one refrigerator would clearly match a sentence referring to a refrigerator. On the other hand, the ReferIt3D dataset classifies “hard” contexts as those with more than two distractor objects, and identifies “view-dependent” instances where recognizing the target requires the observer to adopt a specific viewpoint within the scene. Varied methodologies are employed for assessment; some compute the average IoU [81, 50], whereas others ascertain the average distance between bounding box centroids [81]. In the context of multi-object grounding, the F1 score serves as the metric [82]. Initially, a one-to-one correspondence is established between predicted and actual bounding boxes based on IoUs. Subsequently, pairs surpassing a specified IoU threshold are classified as true positives.
@@ 3.2 Evaluation Metrics
4 Methods
Most of the existing T-3DVG methods are proposed with the fully-supervised learning paradigm. Early approaches primarily utilized pre-trained object detectors or segmentation models to pre-sample all potential objects from the complicated 3D scene. Then, these object proposals are then paired with a given sentence query to learn their semantic correspondence through cross-modal interaction to determine the most matching one. However, this two-stage framework not only severely relies on the quality of the 3D detectors and segmentation models, but also requires time-consuming reasoning among the multiple proposals. To alleviate this limitation, without using object proposals, the one-stage framework is proposed to take the whole 3D scene as the input and embed point-level features in an end-to-end manner. After point-text cross-modal reasoning, it directly regresses the spatial bounding box of the text-guided object. Although the above two types of fully-supervised methods have achieved significant performance in recent years, they require a large number of bounding boxes for numerous objects in the 3D point cloud scenes, along with their descriptive texts, to provide reliable supervision. Considering manually annotating these object-text pairs is very time-consuming and labor-intensive, some works propose to solve the T-3DVG task in the weakly-supervised setting, which only utilizes text labels for grounding without relying on any object bounding box annotations. There are also some recent works that utilize other-modal input (instead of point cloud) for grounding or incorporate traditional grounding models with large language models.
In this section, we provide a complete review of existing T-3DVG methods via a taxonomy in Fig. 10, based on the methods’ architectures and learning algorithms, to better categorize TSGV approaches. Specifically, we discuss the unique characteristics of each category and illustrate the highlights of each T-3DVG method in details.
4.1 Fully-Supervised Methods
4.1.1 Two-Stage Methods
Most existing T-3DVG methods [7, 8, 65, 50, 27, 28, 47, 21, 22, 23, 55, 83, 56, 84, 66, 85, 86, 58, 82, 63, 87, 88, 89, 90, 29, 57, 91, 92, 93] are deployed in a two-stage framework, which first utilizes a pre-trained 3D detection or segmentation model to obtain all possible 3D objects in the scene and then individually matches and reasons them with the query for selecting the best one. These works can be further categorized into a basic proposal-then-selection approach and a multi-modal enhanced proposal-then-selection approach. In particular, the latter approach integrates more contextual modalities’ information into the 3D reasoning process for enriching the representation learning.
Basic proposal-then-selection methods. ScanRefer [7] (as shown in Fig. 12) and ReferIt3D [8] are the pioneer methods proposed to address the T-3DVG task following a basic proposal-then-selection methods. They generally first utilize a pre-trained 3D object detector to obtain all possible objects within the whole 3D scene. Then, they match the objects’ features with the semantic of text descriptions for determining the best aligned object as the final output. Instead of using pre-trained 3D objectors, a few works [65, 50] also introduce to utilize a pre-trained segmentor to obtain fine-grained semantic features for text matching. Considering the object detector is relatively lightweight and enough to capture object-level features, almost all two-stage methods follow the detection-based proposal-then-selection framework.
Figure 12: ScanRefer architecture, reproduced from Chen et al. [7].
Besides, there are also some works [21, 22, 23, 84, 58, 82, 89] designed to address the task-specific challenges in T-3DVG. Considering that object locations are important to reason the spatial relations mentioned in the sentence, spatial embeddings [21, 22] are designed to embed positional-encoded bounding box information into the attention mechanism. To help distinguish complicated background objects from the foreground target, He et al. [23] propose to separately align low-level and high-level of words with the objects for comprehending and recognizing the corresponding object entity and relation. Kato et al. [84] introduce the ARKitSceneRefer method to tackle the object location on small indoor 3D objects. It introduces a new small-object dataset and develops a simple coarse-to-fine framework to achieve the goal. Hsu et al. [58] propose to translate language into programs with hierarchical structures by leveraging large language-to-code models and extend prior neuro-symbolic visual reasoning methods by introducing functional modules that effectively reason about high-arity relations (i.e., relations among more than two objects), key in disambiguating objects in complex 3D scenes. This framework can alleviate two issues: the expense of labeling and the complexity of 3D grounded language. Zhang et al. [82] introduce a new case when multiple objects exist in the same scene, and devise a contrastive learning strategy to distinguish the objects for better determining. Considering unstructured natural utterances and scattered objects might lead to undesirable performances, Wu et al. [89] propose to leverage language models to parse language description, suggesting a referential order of anchor objects. Such ordered anchor objects allow the model to update visual features and locate the target object during the grounding process.
Moreover, some proposal-then-selection works [55, 83, 56, 66, 63, 87] design novel training strategies or relevant joint tasks to assist 3D grounding. For example, Cai et al. [55] and Chen et al. [87] propose to jointly train a 3D captioner and a 3D grounder to provide diverse query descriptions for better supervising the grounding module with mutual information learning. Based on this joint framework, Chen et al. [83] further incorporate a teacher-student architecture into the model design to train a robust grounder by mimicking the strong ability of the joint framework. Zhu et al. [56] propose a pre-training strategy with transformers for 3D vision and text alignment, which utilizes self-attention layers for both single-modal modeling and multi-modal fusion without any sophisticated task-specific design. Specifically, it collects diverse prompts, scene graphs, 3D scans, and objects with self-supervised pre-training. Jin et al. [66] also propose a 3D vision-language pertaining strategy with context-aware spatial-semantic alignment and mutual 3D-language masked modeling to learn more representative features. It shows great performance on the 3D grounding task. Unal et al. [63] develop three novel stand-alone modules to improve grounding performance for challenging repetitive instances, i.e., instances with distractors of the same semantic class. Two new datasets [95, 96] are proposed to scale the knowledge learning-ability of T-3DVG.
Multi-modal enhanced proposal-then-selection methods. Although the existing basic proposal-then-selection methods have significant performance, they solely exploit the contexts within the original 3D point cloud input for retrieving the specific objects according to different queries, lacking sufficient yet complementary information from other views or modalities for enriching the retrieval knowledge. To this end, there are many recent two-stage T-3DVG methods [29, 57, 91, 92, 97, 98, 30, 44, 93] proposed to perceive more contextual information from other perspectives (such as 2D images, multi-view contexts, etc.) for enriching the traditional 3D grounding knowledge, as shown in Fig. 13.
Considering the traditional point cloud data are sparse, noisy, and contain limited semantic information compared with 2D images, many works [29, 57, 91, 92, 93] propose to utilize additional 2D image input to assist 3D visual grounding. These methods generally first detect potential object proposals through 3D object detector, then encode each object’s 2D image into feature vectors to combine with 3D object features for aligning with the semantics of sentence description. Since more information is utilized, these works achieve better performance than basic proposal-then-selection frameworks. In addition to the point cloud input, Xiao et al. [97] propose to build a new representation for 3D points in the semantic point cloud defined on high-level semantics. This new representation guides the model to understand more relationships by simplified representation of objects, rather than limited to traditional appearance learning. Both semantic and appearance object features are concatenated to feed into the latter cross-modal reasoning module. To make T-3DVG more practical in real-world applications, Lu et al. [98] introduce a human-robot interaction fashion that takes human guidance as a cue to facilitate the development of 3D visual grounding. A point cloud of a human agent is specially introduced as additional input along with a designed gesture extraction module. Therefore, the object-level scene representations are reasoned with both human guidance and textual semantics to determine the location of target object. Besides, multi-view learning designs [30, 44] is also a promising multi-modal enhancement strategy. They take multi-view point clouds of the same 3D scene as inputs, and mine their correspondences between the same object for grasping and comprehending the view knowledge. Overall, these multi-modal enhanced two-stage methods achieve better performance as more contextual information is embedded into the representation learning.
Figure 14: Uni3DL architecture, reproduced from Li et al. [99].
4.1.2 One-Stage Methods
Without relying on the quality of pre-trained object generators (i.e., 3D detector or segmentor), some recent T-3DVG methods follow the one-stage framework that trains the 3D grounding framework end-to-end from the feature extraction to the final cross-modal grounding. Among them, inspired by the already successful two-stage approach, some works [24, 100, 99, 101, 102, 103, 104, 70, 71] propose a unified detection-interaction method that jointly trains both object detection and object reasoning with mutual learning. To reduce the training parameters and resources, recent works [25, 26, 67, 105, 69] directly encode the point-wise features of the whole point cloud and globally match all points with the query to regress the optimal locations of object boundary box locations.
Figure 15: 3D-SPS architecture, reproduced from Luo et al. [25].
Based on the powerful grounding ability of 3D-SPS and EDA, the latter works [67, 105] focus more on the detailed network designs. Wang et al. and Jain et al. [67, 69] emphasize the importance of relative spatial relationships among objects described by the sentence description, and propose a 3D relative position module to learn and reason the spatial relations between the key point features before regressing their corresponding bbox outputs. Huang et al. [24] develop a contextual global transformer with learnable queries to iteratively generate text-described objects in the 3D scene. Then, it develops a fine-grained local transformer to associate the object features for scoring and selection. This paper designs more efficient detection module within the grounding pipeline, achieving end-to-end training and reducing the resource costs. Moreover, unlike previous pre-set objects containing lots of background objects, the generated objects of this paper are text-related, significantly improving the bbox quality. Lin et al. [105] incorporate the superpoint mask segmentation into the object grounding task with a mutual learning framework for joint training, bring additional mask-level knowledge for improving the grounding performance. Compared to unified detection-interaction methods, these regression-based methods are more efficient.
Figure 16: WS-3DVG architecture, reproduced from Wang et al. [31].
4.2 Weakly-Supervised Methods
Want et al. [31] propose the first weakly-supervised framework for T-3DVG task, as shown in Fig. 16. To fill the information gap between 3D visual features and textual features, it first extracts all possible objects from the whole 3D scene with a pre-trained 3D detector. Then, it coarsely selects the top-$K$ object proposals based on their semantic similarities with the sentence description. To further mine their potential relationships, it borrows the text-based pre-training strategies from transformer [52], which masks crucial words in the sentence and reconstructs the masked keywords using each selected proposal one by one. The reconstructed accuracy finely reflects the semantic similarity of each object to the query. Therefore, the best matched object is chosen as the final grounding result. Xu et al. [106] also introduce a 3D-VLA network to address the weakly-supervised T-3DVG task. It leverages additional 2D semantics to assist the grounding process. Specifically, it encodes object-level features with pre-trained 3D detector and image-text features with CLIP [61]. Then, it interacts 3D-level, image-level, and text-level semantics to determine which object is matched with the textual description. In summary, there are a few works that make attempts to investigate the challenging weakly-supervised T-3DVG problem. However, their frameworks are still limited without special designs of robust representation learning, resulting in worse performance than existing fully-supervised methods.
4.3 Methods with Other-Modality Input
Instead of using point cloud as the 3D grounding input, there are also a few works [62, 107, 108] proposed to address the task from a new perspective with other types of inputs, such as RGB-D images and human-interaction vectors. Liu et al. [62] propose to ground objects in single-view RGBD images. Given an RGBD image and a sentence query input, they first fuse the multi-modal features via attention mechanisms and generate a visual heatmap to coarsely localize the relevant regions in the RGBD image. Then, an adaptive feature learning based on the heatmap is introduced to perform the object-level matching to finely locate the referred object. Different from previous point-cloud-based grounding, this RGBD-based grounding requires fewer resources to collect and annotate data. Another RGBD-based method [107] is also proposed to handle the holistic multi-modal 3D perception suite. In addition to the RGB-D input, this method also feeds corresponding RGB images into the model. The whole network encodes both 2D and 3D visual features to densely interact with sentence queries, leading to more fine-grained object grounding. Instead of relying on RGB-D data input, Zhan et al. [108] propose to ground target objects in monocular images. As we know, the task of monocular object detection [109, 110] task is well developed and has a large amount of collected data, providing more robust training. Besides, PATRON [102] takes multiple views of non-verbal interactions to locate the target objects in more practical robotics scenarios. Lin et al. [111] also introduce a wild case of automatic driving data and corresponding 2D images for grounding. Overall, the above methods provide a new perspective for T-3DVG instead of understanding and reasoning point clouds.
Figure 17: LAMM architecture, reproduced from Yin et al. [100].
4.4 Methods using Large Models
By benefiting from the strong comprehension of large language models (LLMs), there are many recent T-3DVG methods [112, 108, 44, 104, 113, 85, 89, 100, 114, 96, 115] that adopt large models to construct their grounding framework. Considering the description annotations are complicated but with limited information, most works [112, 108, 44, 104] utilize LLM tools [116, 117, 118, 119, 115] to decompose complex natural language queries into semantic constituents for adaptive grounding, and enrich the contexts of original description with designed prompts for providing more detailed commonsense knowledge, or directly generate a new description from another perspective via captioners. Yang et al. [112] further utilize LLMs to evaluate the spatial relations among the proposed objects to make a final grounding decision. These methods incorporated more diverse textual semantics into the text-object reasoning process, leading to more precise grounding results. Besides, some works [113, 85, 89] try to leverage the commonsense reasoning capability of LLMs to make inferences about the objects in the scene and their interactions. Specifically, Bakr et al. [85] and Wu et al. [89] propose to exploit Chain-of-Thoughts [120, 121] to decompose the referring task into multiple interpretable steps, whereas to reach the final target the model must first predict the anchors one by one, in a logical order. However, the above methods simply take the LLMs as a module tool to assist the traditional grounding model, failing to exploit the strong power of the LLM architecture designs. To this end, recent methods [100, 114, 96] extend the reasoning ability of multi-modal LLMs [122, 123, 124] into the 3D scenarios, as shown in Fig. 17. Specifically, they design object or scene-level 3D encoders to embed the visual input into the same latent vectors as the language dimensions in LLMs. Then, both textual embeddings and visual embeddings are jointly fed into the LLMs for reasoning. The grounding results can be achieved by inputting a specific prompt into the LLMs.
@@ 4.4 Methods using Large Models
TABLE III: Performance Comparison on ScanRefer dataset.
@@
4.5 Discussion
In summary, existing T-3DVG methods can be generally categorized into two types: fully-supervised approaches and weakly-supervised approaches. As for fully-supervised approaches, most works follow a basic proposal-then-selection framework, which first utilizes pre-trained models to extract all possible object proposals in each scene and then match them with textual sentences for selection. To enhance the prior knowledge for assisting grounding, some proposal-then-selection methods also utilize multi-modal data as additional input to provide sufficient grounding information. Although these two-stage framework achieves significant performance, they severely rely on the quality of the pre-trained model and costs large resources to compute backgrounding objects. Therefore, one-stage methods are proposed to directly regress the target objects. Some of them follow the two-stage methods to design a unified detection-interaction frameworks, which replace the pre-trained object model with an efficient learnable one. The others directly regress the target object with point-level feature reasoning. As for weakly-supervised approaches, they solely exploit the knowledge of the relationship annotations between scenes and queries without object-level bounding boxes. Existing weakly-supervised works generally follow a proposal-then-matching framework that scores potential proposals with self-learning strategies. There are also some works that utilize other 3D-modal input instead of point cloud, or combine LLMs with traditional grounding models. Overall, there is still a lot of room for exploration in researching methods of weak-supervised approaches and methods of using large models.

### v2

@@ IV-A Benchmark Datasets
Figure 16: Demonstration of the statistics of 18 general categories, consisting of 256 sub-categories within a 3D indoor environment, predominantly featuring furniture items like chairs, tables, and cabinets. The terms ”Maximum,” ”Average,” and ”Minimum” signify the highest, mean, and lowest counts, respectively, for a specific type of object present in a scene. For instance, a scene may contain as many as 32 chairs.
@@
TABLE III: Performance Comparison on ScanRefer dataset. ’*’ denotes training with few-shot annotations (i.e., 10% data).
@@
G3-LQ [108] CVPR 2024 fully one-stage 3D 88.59 73.28 50.23 39.72 55.95 44.72
MA2TransVG [90] CVPR 2024 fully two-stage 3D 86.3 74.1 53.8 41.4 57.9 45.7
VPP-Net [91] CVPR 2024 fully two-stage 3D 86.05 67.09 50.32 39.03 55.65 43.29
MCLN [109] ECCV 2024 fully one-stage 3D 84.43 68.36 49.72 38.41 54.30 42.64
3D-CTTSF [94]* ACMMM 2024 fully two-stage 3D 76.75 58.27 38.60 28.98 46.00 34.67
@@
3D-CTTSF [94]* ACMMM 2024 fully two-stage 3D+2D 82.87 61.79 39.57 28.89 47.97 35.28
IV-B Evaluation Metrics
In the realm of 3D grounding assessment, there is a necessity to compare predicted locations, typically depicted as bounding boxes, against ground truth locations of objects within testing samples. Acc@$K$IoU [11] stands as the predominant metric in 3D visual grounding, quantifying the proportion of positive predictions surpassing a designated intersection over union (IoU) threshold, typically denoted as $K$ and commonly set at 0.25 or 0.5. It is notable that certain datasets present varying evaluation configurations. For instance, in datasets like Nr3D and Sr3D, predefined object proposals are provided, with the metric focusing on the precision of selecting the target bounding box from these proposals. Additionally, various datasets assess performance across diverse scenarios. ScanRefer [11] partitions datasets into unique/multiple/overall splits, while ReferIt3D [12] categorizes them into easy/hard/view-dependent/view-independent/overall divisions. Specifically, in the ScanRefer dataset, the “unique” split includes instances where only a single distinct object from a specific category fits the description, while the “multiple” split involves cases with several objects from the same category, creating ambiguity. For example, a scene with only one refrigerator would clearly match a sentence referring to a refrigerator. On the other hand, the ReferIt3D dataset classifies “hard” contexts as those with more than two distractor objects, and identifies “view-dependent” instances where recognizing the target requires the observer to adopt a specific viewpoint within the scene. Varied methodologies are employed for assessment; some compute the average IoU [142, 56], whereas others ascertain the average distance between bounding box centroids [142].
@@ IV-B Evaluation Metrics
G3-LQ [108] CVPR 2024 fully one-stage 3D - 66.3 - - 73.1
MA2TransVG [90] CVPR 2024 fully two-stage 3D 76.0 69.3 64.5 73.8 73.9
VPP-Net [91] CVPR 2024 fully two-stage 3D - - - - 68.7
MiKASA [86] CVPR 2024 fully two-stage 3D 78.6 67.3 70.4 75.4 75.2
MCLN [109] ECCV 2024 fully one-stage 3D - - - - 68.4
GNL3D [95] ACMMM 2024 fully two-stage 3D 72.8 64.0 58.0 70.6 70.1
@@
G3-LQ [108] CVPR 2024 fully one-stage 3D - 50.7 - - 58.4
MA2TransVG [90] CVPR 2024 fully two-stage 3D 71.1 57.6 62.5 65.4 65.2
VPP-Net [91] CVPR 2024 fully two-stage 3D - - - - 56.9
MiKASA [86] CVPR 2024 fully two-stage 3D 69.7 59.4 65.4 64.0 64.4
MCLN [109] ECCV 2024 fully one-stage 3D - - - - 59.8
@@
V Performance Comparison
We have summarized the performance of various already-published T-3DVG methods on the ScanRefer, Sr3D, and Nr3D benchmark datasets since 2021 in Tables III, IV, and V in chronological order, respectively. In these tables, we also list the framework type (two-stage or one-stage), the input modality, and the supervision method (fully-supervised or weakly-supervised) for each method. Overall, two-stage methods tend to perform slightly worse than one-stage methods in grounding due to their high dependency on the performance of pretrained 3D detection models, which can overlook unimportant or inconspicuous objects. However, some recent two-stage methods [81, 28] have addressed this shortcoming by refining the detected object locations through text guidance or other strategies. For the same method, results are always better when the input includes additional multiview 2D information compared to those using only the 3D modality as input. In the tables, it can also be observed that methods employing multi-task joint training [61, 83, 62] often achieve relatively better results. This improvement can be attributed to the benefits of more data and the learning of more comprehensive feature representations brought by learning from different tasks.
By comparing the results across three datasets, it can be observed that the relative performance of methods is generally consistent across different datasets. However, there are subtle differences between the tasks of the ScanRefer dataset and the two datasets of Referit3D (Sr3D, Nr3D). ScanRefer dataset requires methods to regress the bounding box of the target object, while Referit3D datasets only require selecting among the provided bounding boxes. Therefore, some one-stage methods have a more significant advantage on ScanRefer due to their ability to achieve higher quality bounding boxes, whereas this advantage is less noticeable on Sr3D and Nr3D dataset. The comparison between EDA [32] and ViL3DRel [28] provides a clear example of this. Note that the results of the CORE-3DVG [81] on the Referit3D dataset are based on the detected bounding boxes for grounding, rather than the conventional ground truth bounding boxes, resulting in slightly inferior performance compared to the results on the ScanRefer.
Additionally, to ensure a fair comparison, we have listed all methods involving large language models (LLMs) in another Table VI, regardless of whether LLMs are used for data pre-processing or direct training of multi-modal LLMs. As can be seen from the table, due to the varying use of LLMs, the results of these methods differ significantly. Methods that employ instruction-tuning with multimodal LLMs [101] perform worse compared to common T-3DVG methods; conversely, using LLMs for commonsense reasoning on processed information [117] yields impressive results. This variance in performance underscores the complexity and potential of integrating LLMs in T-3DVG tasks.